# Zero-free regions for the independence polynomial on restricted graph classes

### v1

@@ 6. The case when $H$ is neither a subdivided claw nor a path
Given a graph $H$ which is neither a path nor a subdivided claw, then either $H$ has a vertex of degree at least $4$ or $H$ has (at least) two vertices of degree $3$. In the latter case, set $k\in\mathbb{N}$ to be larger than the distance between these vertices of degree $3$ and and in the former case, set $k=0$. Let $\mathcal{T}=\{T_{d}:d\in\mathbb{N}\}$ be the family of trees, where $T_{d}$ is obtained from the complete binary tree of height $d$ by subdividing each edge $2k$ times (so there are $2k$ intermediate vertices between branch vertices). Refer to Figure 3. By our choice of $k$, all trees in $\mathcal{T}$ are $H$-free.
@@
At a fixed point, $g_{\lambda}(z)=z$ and $z=(1\pm o(1))\lambda^{1/3}$. Letting $\sim$ denote equality up to a factor $(1\pm o_{\lambda}(1))$, and $z=z_{\lambda}$ the identified fixed point, we have $z_{\lambda}\sim\lambda^{1/3}$. Also,
@@
We now employ [11, Lemma 13]. In our application, $\Delta=2$, and $H_{\Delta}=H_{2}$ is the set of all functions that can be obtained by composing sequences composed of $f_{\lambda,1}$ and $f_{\lambda,2}$. The lemma asserts that the existence of an indifferent fixed point at $\lambda=\lambda_{0}$ (just established) ensures that $\lambda_{0}$ is an accumulation point of zeros realised by balanced trees with branching factors in $\{1,2\}$. In our application, however, we want to ensure that pairs of degree-2 vertices remain separated, by $2k$ degree-1 vertices. In fact, the trees $T\in H_{2}$ referred to in the statement of Buys’ Lemma 13 do have this property. However, we have to peek into the proof to see this.
The proof of the lemma first establishes that a certain function $g_{\lambda}^{\circ n}\circ h_{\lambda}$ is not ‘normal’ about $\lambda=\lambda_{0}$. The function $g_{\lambda}$ and the point $\lambda_{0}$ are the same as ours, so the issue for us is whether the $h_{\lambda}\in H_{2}$ has two occurrences of $f_{\lambda,2}$ that are too close together. Consulting Remark 8, we discover that $h_{\lambda}$ is in fact an initial segment of the composition $f_{\lambda,2}\circ f_{\lambda,1}^{\circ 2k}$, so in fact contains at most one occurrence of $f_{\lambda,2}$. Finally, at the end of the proof of Lemma 13, either one or two functions are removed from the composition of functions $g_{\lambda}^{\circ n}\circ h_{\lambda}$ to yield the desired function $\tilde{g}_{\lambda}$. Thus, $\tilde{g}_{\lambda}$ is a composition of a contiguous subsequence of functions from
@@ Acknowledgments
We thank Guus Regts and Ferenc Bencs for helpful discussions and feedback on an earlier draft.

### v2

@@ 6. The case when $H$ is neither a subdivided claw nor a path
Given a graph $H$ which is neither a path nor a subdivided claw, then either $H$ has a vertex of degree at least $4$ or $H$ has (at least) two vertices of degree $3$. In the latter case, set $k\in\mathbb{N}$ to be larger than the distance between these vertices of degree $3$ and in the former case, set $k=0$. Let $\mathcal{T}=\{T_{d}:d\in\mathbb{N}\}$ be the family of trees, where $T_{d}$ is obtained from the complete binary tree of height $d$ by subdividing each edge $2k$ times (so there are $2k$ intermediate vertices between branch vertices). Refer to Figure 3. By our choice of $k$, all trees in $\mathcal{T}$ are $H$-free.
@@
With a view to verifying that $g_{\lambda}$ has a fixed point $z_{\lambda}$ close to $\lambda^{1/3}$, let $\varepsilon>0$ and define $z^{\pm}_{\lambda}=(1\pm\varepsilon)\lambda^{1/3}$. Then
[…]
$=\frac{\lambda^{2}(1+O(\lambda^{-2/3}))}{(1+\varepsilon)^{2}\lambda^{5/3}(1+O(\lambda^{-1/3}))}$
$=(1+\varepsilon)^{-2}\lambda^{1/3}(1+O(\lambda^{-1/3})).$
Thus, provided $\lambda\gg\varepsilon^{-3}$, we have $z^{+}_{\lambda}-g_{\lambda}(z_{\lambda}^{+})>0$ and $z^{-}_{\lambda}-g_{\lambda}(z_{\lambda}^{-})<0$. It follows that $g_{\lambda}$ has a fixed point $z_{\lambda}$ satisfying $z_{\lambda}^{-}<z_{\lambda}<z_{\lambda}^{+}$, at which $z_{\lambda}=\lambda^{1/3}(1+O(\lambda^{-1/3}))=\lambda^{1/3}+O(1)$.
Letting $\sim$ denote equality up to a factor $(1\pm o_{\lambda}(1))$,
@@
We now employ [9, Lemma 13]. In our application, $\Delta=2$, and $H_{\Delta}=H_{2}$ is the set of all functions that can be obtained by composing sequences composed of $f_{\lambda,1}$ and $f_{\lambda,2}$. The lemma asserts that the existence of an indifferent fixed point at $\lambda=\lambda_{0}$ (just established) ensures that $\lambda_{0}$ is an accumulation point of zeros realised by balanced trees with branching factors in $\{1,2\}$. In our application, however, we want to ensure that pairs of degree-2 vertices remain separated, by $2k$ degree-1 vertices. In fact, the trees $T\in H_{2}$ referred to in the statement of Buys’ Lemma 13 do have this property. However, we have to peek into the proof to see this.
The proof of the lemma first establishes that a certain function $g_{\lambda}^{\circ n}\circ h_{\lambda}$ is not normal about $\lambda=\lambda_{0}$ (or more precisely that the family $\{\lambda\mapsto g_{\lambda}^{n}\circ h_{\lambda}(0)\}_{n\geq 1}$ is not normal in any neighbourhood of $\lambda_{0}$). The function $g_{\lambda}$ and the point $\lambda_{0}$ are the same as ours, so the issue for us is whether the $h_{\lambda}\in H_{2}$ has two occurrences of $f_{\lambda,2}$ that are too close together. Consulting Remark 8, we discover that $h_{\lambda}$ is in fact an initial segment of the composition $f_{\lambda,2}\circ f_{\lambda,1}^{\circ 2k}$, so in fact contains at most one occurrence of $f_{\lambda,2}$. Finally, at the end of the proof of Lemma 13, either one or two functions are removed from the composition of functions $g_{\lambda}^{\circ n}\circ h_{\lambda}$ to yield the desired function $\tilde{g}_{\lambda}$. Thus, $\tilde{g}_{\lambda}$ is a composition of a contiguous subsequence of functions from
@@ Acknowledgments
We thank Guus Regts and Ferenc Bencs for helpful discussions and feedback on an earlier draft. We also thank the anonymous referee for their careful reading and valuable feedback on this work.